A survey of 500 top Canadian companies reported what the companies' hiring outlook was for the next 18 months as well as the general company outlook for the economy over the same period of time.

Economy Outlook Add Jobs Hiring Outlook No Change Cut Jobs
Favourable 100 50 25
Unknown 50 150 60
Unfavourable 15 10 40



Given that a randomly selected company has a favourable forecast, what is the probability that it plans to cut jobs during the next 18 months?
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

P(a favourable outlook for the economy | cut jobs) = P(a favourable outlook for the economy and cut jobs) / P(favorable forecast)

=(25/500)/(175/500)=25/175